OpenStudy (anonymous):

easy question but just having a brainfart... x^2+x^2 = 2x^2 correct?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

great thanks

OpenStudy (anonymous):

Np

hero (hero):

x^2 + x^2 = (1 + 1)x^2 = 2x^2

OpenStudy (anonymous):

cool. thank you